OpenStudy (goldrush18):

the expression 2x^3+ax^3+bx+6 is divisible by x-2 and gives a remainder of -12 when divided by x+1. find the values of a and b and factorize the expression completely the expression 2x^3+ax^3+bx+6 is divisible by x-2 and gives a remainder of -12 when divided by x+1. find the values of a and b and factorize the expression completely @Mathematics

OpenStudy (anonymous):

replace x by 2, set expression = 0 and get an equation in a and b replace x by -1, set equal to -12 and get another equation in a and b solve the system for a and b

OpenStudy (goldrush18):

what do u get when u do the simultaneous equation

OpenStudy (anonymous):

is there a typo here because you have \[2x^3+ax^3+bx+6\] and i am thinking it should be \[2x^3+ax^2+bx+6\]

OpenStudy (anonymous):

clarify this and i will work it out for you

OpenStudy (goldrush18):

yes typo sorry

OpenStudy (goldrush18):

the second one u wrote is the right one

OpenStudy (anonymous):

ok so replace x by 2 and get \[2\times 2^3+a\times 2^2+b\times 2+6=0\] or \[16+4a+2b+6=0\] or \[4a+2b=-22\]

OpenStudy (anonymous):

then replace x by -1 and get \[2\times (-1)^3+a\times (-1)^2+b\times (-1)+6=-12\] \[-2+a-b+6=-12\] \[a-b=-16\]

OpenStudy (anonymous):

so your two equations are \[4a+2b=-22\] \[a-b=-16\] multiply the second by 2 to get \[4a+2b=-22\] \[2a-2b=-32\] \[6a=-54\] \[a=-9\]

OpenStudy (anonymous):

then \[-9-b=-16\] \[-b=-7\] \[b=7\]

OpenStudy (anonymous):

check my arithmetic, because i just typed it and didn't do it with pencil and paper, but it looks right

OpenStudy (goldrush18):

ok im checking i have more questions to post hoping u can help me along the way